OpenStudy (anonymous):

What is the measure of one angle of a regular 16-gon? If necessary, round to the nearest degree. 120 135 140 158

OpenStudy (anonymous):

180 - (360/ 16) = 180 - 22 =158

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw